Let the random variable x represent the number of cars owned by a family. Assume that...

  1. Let the random variable x represent the number of cars owned by a family. Assume that x can take on five values: 0, 1, 2, 3, 4. A partial probability distribution is shown below:

        

x

0

1

2

3

4

p(x)

0.2

0.1

0.3

?

0.1

         i.       The probability that a family owns three cars equals _______

         ii.       The probability that a family owns between 1 and 3 cars, inclusive, equals _______

         iii.      The expected number of cars equals _______
Answer #1

i) Since  Σρ(α) = 1 T=0

0.2+0.1 +0.3+p(3) +0.1 = 1

+0.7+p(3) = 1

→p(3) = 1 -0.7 = 0.3

ii) The probability that a family owns between 1 and 3 cars, inclusive, equals

=p(1) +p(2) +p(3)

= 0.1 +0.3 +0.3

= 0.7

iii) The expected number of cars equals

E(X)= x*p(2) = (0*0.2)+(1+0.1)+(2+0.3)+(3*0.3)+(4*0.1) = 2 r=0